Please help me with this word problem: A 24-foot ladder resting against a house reaches a windowsill 16 feet above the ground. How far is the foot of the ladder from the foundation of the house? Express your answer to the nearest tenth of a foot.
Answer by stanbon(75887)

A 24-foot ladder resting against a house reaches a windowsill 16 feet above the ground. How far is the foot of the ladder from the foundation of the house? Express your answer to the nearest tenth of a foot.
-------------------
Draw the picture of the ladder leaning against the house.
You have a right triangle with hypotenuse = 24
and height = 16.
Find the base using Pythagoras:
----
base^2 = hypotenuse^2 - height^2
----
b^2 = 24^2-16^2
b^2 = 320
b = sqrt(320)
b = 17.9 ft
================
Cheers,
Stan H.